OpenStudy (anonymous):

An arc length 5 feet subtends central thetaa,in a circle with a piameter 7 feet. a)Determine the area of the exact radian measure of theta b)Determine the area of the sector with central angle theta c)determine the peremeter and the sector

OpenStudy (mertsj):

What is the circumference of the circle?

OpenStudy (anonymous):

360 degrees

OpenStudy (mertsj):

Circumference of a circle is pi times diameter. What is the circumference of the circle?

OpenStudy (anonymous):

pie r ^2?

OpenStudy (mertsj):

\[C=\pi d\]

OpenStudy (anonymous):

is d diameter?

OpenStudy (mertsj):

yes. Didn't you read what I wrote earlier: Circumference of a circle is pi times diameter.

OpenStudy (anonymous):

yea i am just making sure

OpenStudy (anonymous):

so c would be pie7

OpenStudy (mertsj):

Yes. The circumference is 7 pi.

OpenStudy (mertsj):

What is the measure of the arc according to the problem?

OpenStudy (anonymous):

the arc is 5 feet

OpenStudy (mertsj):

So the arc is this fractional part of the circle: \[\frac{5}{7\pi}\]

OpenStudy (mertsj):

Do you understand that?

OpenStudy (anonymous):

yes i do

OpenStudy (mertsj):

I see that you know a circle is 360 degrees. How many radians is a circle?

OpenStudy (anonymous):

ummmm im not sure is it 4

OpenStudy (mertsj):

360 degrees = 2 pi radians

OpenStudy (anonymous):

aw k i didnt know

OpenStudy (mertsj):

So a circle is 2 pi radians. So find theta. It is: \[\frac{5}{7\pi} \times 2\pi \]

OpenStudy (anonymous):

it =10

OpenStudy (mertsj):